# A machine-learning photometric classifier for massive stars in nearby galaxies

### v1

@@ 4.3 RSG luminosity functions
Figure 5: Luminosity functions for all sources identified as RSGs in M31, M33 and NGC 6822. We note the presence of some very luminous sources for M31 and M33 with $\textrm{log}(L/L_{\odot})\,$¿5.5, indicated by a dashed line (see Section 4.3).
@@
Given that the highest success rate is obtained for RSGs we opted to explore their luminosity functions for the galaxies with sufficient numbers, namely M31, M33, and NGC 6822. Although historically the Humphreys-Davidson limit (Humphreys & Davidson, 1979) is found at $\textrm{log}(L/L_{\odot})\,$$\sim 5.8$ dex, more recent works lowered the limit to $\textrm{log}(L/L_{\odot})\,$$\sim 5.5$ dex (Davies et al. 2018 for the LMC and SMC; McDonald et al. 2022 for M31). Hence, it is very intriguing to investigate predictions of RSGs above this limit.
@@
For this, we need first to estimate the luminosity for each source. We employed the bolometric correction for RSGs from Neugent et al. (2020) following their section 5, given the $K_{s}-\textrm{magnitude}$ as provided by the UHS survey (using a 3” aperture for $JK_{s}$ photometry). To correct for extinction, we assumed a uniform value of $A_{V}=0.75$ mag as they did, but for all sources. When not applying additional extinction for the brighter sources, our estimated luminosities are similar to those from (Wang et al., 2021), which are calculated by integrating the spectral energy distributions, a more robust method that takes into account the emission from the dust shell. For NGC 6822, we assumed $E(B-V)=0.25$ mag (Massey et al., 2007b). The luminosities of RSGs in NGC 6822 are underestimated by 0.05 dex compared to the luminosities of the RSGs in common with Antoniadis et al. (2025), which were calculated by integrating the spectral energy distributions. The results are presented in Fig. 5 (we note here that the uncertainty in this estimate due to the correction is 0.05 dex; Neugent et al. 2020).
@@
From this plot, it is striking that there are several luminous sources above $5.5\,L_{\odot}$. In particular we find 22 M31 and 6 M33 sources. In Table 5 we present them by including their Spitzer and Gaia DR3 IDs, their Gaia parallax and proper motions (whenever available), previous (original) and final (predicted) classifications, as well as their final probability, band completeness, and estimated luminosity.
@@
In M31, the three most luminous sources (close to $\textrm{log}(L/L_{\odot})\,$$\sim 7$; IDs: M31-439614, M31-439351, and M31-439254) do not have a Gaia counterpart, which means that these could be foreground red stars. However, the next most luminous source (ID M31-350) is a confirmed M1 I RSG (J004428.48+415130.9), with a range of $\textrm{log}(L/L_{\odot})\,$$=5.43-5.64$ McDonald et al. (2022). Such luminous sources may actually host a dense and complex circumstellar environment (similar to what has been observed in WOH G64; Ohnaka et al. 2024; Munoz-Sanchez et al. 2024a) that can contribute up to 0.3 dex in luminosity. If we consider these uncertainties in our approach, then our estimate of $\textrm{log}(L/L_{\odot})\,$$=5.9$ could be corrected by 0.35 dex, making it consistent with values from the literature. Based on our Gaia foreground cleaning method, the parallax and proper motion measurements make it a strong M31 candidate. Therefore, it certainly provides an upper bound for the Humphreys-Davidson limit in M31. Below that there are an additional 12 sources (between $\textrm{log}(L/L_{\odot})\,$$\sim 5.8-5.5$), of which 3 have Gaia parallax and proper motion values (making them best candidates of high luminous RSGs [^13] Among these sources is J003951.33+405303.7 from McDonald et al. (2022), corresponding to M31-439741, which we predicted robustly as RSG. Sources J004731.12+422749.1 (M31-1011) and J004539.99+415404.1 (M31-336) were also predicted as RSG, but we were unable to estimate their luminosity. For sources J004428.12+415502.9 and J004520.67+414717.3 we were unable to find a good match within a few arcsec in our catalog, probably due to our different Gaia cleaning approach., another set of 8 sources with Gaia photometry only (with at least one band measurement), and one source without any Gaia match (hence, more probable to be a foreground source).
@@
In M33, we find two sources (M33-179 and M33-520) at $\textrm{log}(L/L_{\odot})\,$$\sim 6$). Since both of them lack Gaia parallax and proper motion measurements (but have photometric measurements) we cannot determine if they are genuine M33 members or not. These are followed by source M33-173 at $\textrm{log}(L/L_{\odot})\,$$\sim 5.7$, for which full Gaia information exists. However, this is a known OB star (Massey et al., 2016) and marks an erroneous prediction as a RSG. Another source (ID M33-646) has only Gaia photometry and, therefore, we can only tentatively consider it as a M33 RSG. The remaining two sources have full Gaia information, and have been classified spectroscopically as RSGs. Therefore, they are the sources with the most robust data regarding their M31 membership status and spectral classification, and their derived luminosity sets a limit at $\textrm{log}(L/L_{\odot})\,$$\sim 6$ (which could be lowered to $\textrm{log}(L/L_{\odot})\,$$\sim 5.7$ if we consider the correction due to the circumstellar contribution).
This exploration emphasizes the need for further detailed investigations into these luminous RSGs, to confirm their true nature as RSGs, and to accurately determine their luminosities through spectral energy distribution fitting. Precise luminosity measurements will also aid in exploring the Humphreys-Davidson limit (Humphreys & Davidson, 1979). JWST is the perfect instrument to provide images that will decisively determine (if and) how much these sources are contaminated.
@@ 4.4 Dusty YSGs
Figure 6: IR color ($[3.6]-[4.5]$) distribution of all (best candidate) sources predicted as YSGs. The violet curve corresponds to a Gaussian fit and the black dashed line indicates the $1\sigma$ threshold (0.24 mag) above which we identify sources as dusty YSGs.
@@
SextansA-2751 NGC2403-11707 M31-315890 IC1613-2519 IC1613-27853 NGC6822-19461
M81-12028 M83-5253 M31-318902 IC1613-3532 IC1613-27906 NGC6822-19716
NGC253-1672 M83-7102 M31-373976 IC1613-4592 IC1613-28726 NGC6822-20136
NGC253-1873 M83-13705 M31-389826 IC1613-5907 IC1613-29826 NGC6822-20875
NGC253-2299 M83-22806 M31-408762 IC1613-6759 IC1613-30616 NGC6822-20905
NGC253-3168 Pegasus-11040 M31-439225 IC1613-7221 IC1613-31585 NGC6822-21661
NGC253-3657 Pegasus-15252 M31-439242 IC1613-8299 WLM-89 NGC6822-21893
M33-57845 Pegasus-15778 M31-439244 IC1613-8407 WLM-234 NGC6822-22456
M33-58478 M31-5753 M31-445810 IC1613-8919 WLM-252 NGC6822-23887
M33-59527 M31-5947 M31-445852 IC1613-10537 WLM-13718 NGC6822-24378
M33-61917 M31-16636 M31-487362 IC1613-12631 WLM-16056 NGC6822-24663
M33-66957 M31-53947 M31-510271 IC1613-12668 WLM-19863 NGC6822-24791
M33-68837 M31-134734 M31-517843 IC1613-13160 NGC3109-3334 NGC6822-24814
M33-69046 M31-145674 M31-523223 IC1613-13595 NGC3109-4208 NGC6822-25657
M33-70621 M31-146048 M31-625493 IC1613-13656 NGC3109-8114 NGC6822-27107
M33-72475 M31-148068 M31-762645 IC1613-14055 NGC6822-3644 NGC6822-27519
M33-72577 M31-181722 M31-798387 IC1613-15060 NGC6822-5818 NGC6822-28156
M33-72755 M31-192486 M31-825408 IC1613-17477 NGC6822-6615 NGC6822-28699
M33-75218 M31-234499 M31-840950 IC1613-20305 NGC6822-9575 NGC6822-29242
M33-75973 M31-238753 SextansB-5786 IC1613-20366 NGC6822-9663 NGC6822-30436
M33-78021 M31-249436 SextansB-6281 IC1613-21316 NGC6822-9932 IC10-5584
M33-78763 M31-251868 SextansB-6394 IC1613-24234 NGC6822-11910 IC10-6521
NGC4826-59 M31-269242 SextansB-12820 IC1613-24827 NGC6822-15396 IC10-9582
NGC4826-160 M31-279262 NGC4736-1029 IC1613-25189 NGC6822-16091 IC10-43559
NGC4826-4047 M31-284554 NGC4736-5124 IC1613-26535 NGC6822-18630
NGC2403-4500 M31-295938 IC1613-251 IC1613-26813 NGC6822-19248
NGC2403-5990 M31-297792 IC1613-1015 IC1613-27793 NGC6822-19272
Since our catalog is based on IR photometry, we are particularly sensitive to dusty, evolved stars. This makes us especially effective at detecting and characterizing the population of dusty YSGs, which are likely the evolutionary descendants of luminous RSGs. The ‘RSG problem’, i.e. the observed scarcity of luminous RSGs exploding as Type-II supernovae (e.g. Smartt et al. 2009; Smartt 2015; Davies & Beasor 2020) suggests that the most luminous RSGs may evolve to warmer temperatures, either by undergoing a blue loop phase (e.g. Yoon & Cantiello 2010; Koumpia et al. 2020; Yang et al. 2023; Zapartas et al. 2024) or due to binary interaction because RSG winds are not strong enough to strip RSGs (e.g. Beasor et al. 2020; Decin et al. 2024; Antoniadis et al. 2024). Due to their high mass-loss rates during the RSG phase, these YSGs are often surrounded by significant amounts of material where dust actively forms (Gordon & Humphreys, 2019; Antoniadis et al., 2024; Decin et al., 2024).
To investigate potentially interesting candidates in our sample we considered only the YSGs fulfilling the quality criteria (as set for final probability and band completeness in Section 2.4) in all galaxies. In Fig. 6 we show the distribution of their IR color $[3.6]-[4.5]$. We fit this distribution with a Gaussian function to find a mean value of --0.07 mag (consistent with 0, which is the expected mean value) and a standard deviation of 0.24 mag. This value is also consistent with the value (0.25 mag) that has been used previously (e.g. Kourniotis et al. 2017) based on the LMC and SMC results (see Bonanos et al. 2009, 2010). Hence, we find 159 sources above the $1\sigma$ threshold (indicated in Fig. 6), which we identify as dusty YSGs (corresponding to a fraction of $\sim 0.07$ with respect to their total number of 2071 sources). It is interesting to note that there are 8 sources above the $2\sigma$ value (0.48 mag), namely M33-59527, M33-61917, M33-75218, M31-439225, M31-439242, M31-523223, WLM-19863, and NGC6822-30436), and another two (IC1613-14055 and WLM-234) above the $3\sigma$ threshold (0.72 mag). In Table 6 we present the IDs of all sources above the $1\sigma$ threshold. These are potentially Yellow Hypergiants which have suffered episodic mass loss and can provide a link to the ”RSG problem”. These sources need to be spectroscopically studied to verify their status.
5 Summary and Conclusions
We investigated the effect of metallicity on different stellar populations, finding expected trends, such as a decrease in WR stars at lower metallicities and a relative increase in BSGs and YSGs. However, a number of selection biases, among which Spitzer ’s sensitivity to dusty evolved stars, must be considered when interpreting these trends. We also identified 21 luminous RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 5.5$), including 6 extreme RSGs in M31 ($\textrm{log}(L/L_{\odot})\,$$\geq 6$), challenging the Humphreys-Davidson limit. Further investigation of these sources is necessary to confirm their nature and more accurately determine their luminosity. Additionally, 159 dusty YSGs were detected. These are optimal candidates of Yellow Hypergiants, key sources to understand the ”RSG problem”, and follow-up observations are vital.
Our catalog provides the largest sample of machine-learning-classified massive stars in nearby galaxies, making it a crucial reference for future studies. It enables the identification of prime targets for spectroscopic follow-up, particularly luminous RSGs and YHGs, to further investigate their evolutionary pathways and the role of episodic mass loss in massive star evolution. Furthermore, we compiled a catalog of 5,273 spectroscopically confirmed sources for all the 26 galaxies as derived from the literature. This offers a unique dataset of reference regarding spectral types for all massive stars and candidates known so far (including additional $\sim 330$ other sources). The accuracy of our classifier can be enhanced by incorporating additional spectral data and improving the handling of observational biases. Expanding this method with JWST photometry will allow the study of massive stars in more distant galaxies, providing deeper insights into their role in galactic evolution.
Acknowledgements. GM, AZB, KA, GMS, EC, SdW acknowledge funding support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (“ASSESS”, Grant agreement No. 772086). EZ acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project No: 7933).
Facilities: This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The UHS is a partnership between the UK STFC, The University of Hawaii, The University of Arizona, Lockheed Martin and NASA.
Software: This research made use of NumPy (Harris et al., 2020), pandas (Wes McKinney, 2010; Mckinney, 2011), matplotlib (Hunter, 2007), imbalanced-learn (Lemaître et al., 2017), scikit-learn (Pedregosa et al., 2011), Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013, 2018), Jupyter Notebooks (Kluyver et al., 2016), TOPCAT, an interactive graphical viewer and editor for tabular data (Taylor, 2005).
The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of NASA’s Astrophysics Data System, and SIMBAD database, operated at CDS, Strasbourg, France.
Appendix A Classified sources from the literature

### v2

@@ 4.3 RSG luminosity functions
Figure 5: Luminosity functions for all sources identified as RSGs in M31, M33, and NGC 6822. We note the presence of some very luminous sources for M31 and M33 with $\textrm{log}(L/L_{\odot})\,$¿5.5, indicated by a dashed line (see Sect. 4.3).
@@
Given that the highest success rate is obtained for RSGs, we opted to explore their luminosity functions for the galaxies with sufficient numbers, namely M31, M33, and NGC 6822. Humphreys-Davidson (HD) limit (Humphreys and Davidson, 1979) is a well established empirical upper limit to the luminosity of massive stars. Hot stars are prohibited from moving to cooler temperatures, while LBVs can only sustain short excursions to cooler temperatures, which are considered unstable, and after intense or episodic mass loss they return to their original position in the HRD (Higgins and Vink, 2020; Smith, 2026). These stars do not pass through the RSG phase and they end up as WR stars. For lower mass stars (8-30 M) the evolution leads them all the way up the RSG phase. The HD limit marks the luminosity where substantial mass loss occurs (as they approach their Eddington limit) and is rarely found above that []Davies2018, McDonald2022, Higgins2020. Thus, RSGs either move to hotter positions in the HRD or die as core-collapse supernovae. Although historically the HD limit (Humphreys and Davidson, 1979) is found at $\textrm{log}(L/L_{\odot})\,$$\sim 5.8$ dex, more recent works lowered the limit to $\textrm{log}(L/L_{\odot})\,$$\sim 5.5$ dex (Davies et al. 2018 for the LMC and SMC; McDonald et al. 2022 for M31). Hence, it is very intriguing to investigate predictions of RSGs above this limit.
@@
For this, we first need to estimate the luminosity for each source. We employed the bolometric correction for RSGs from Neugent et al. (2020) following their Sect. 5, given the $K_{s}-\textrm{magnitude}$ as provided by the UHS survey (using a 3” aperture for $JK_{s}$ photometry). To correct for extinction, we assumed a uniform value of $A_{V}=0.75$ mag as they did, but for all sources. When not applying additional extinction for the brighter sources, our estimated luminosities are similar to those from (Wang et al., 2021), which are calculated by integrating the spectral energy distributions, a more robust method that takes into account the emission from the dust shell. For NGC 6822, we assumed E(BV)=0.25E(B-V)=0.25 mag (Massey et al., 2007b). The luminosities of RSGs in NGC 6822 are underestimated by 0.05 dex compared to the luminosities of the RSGs in common with Antoniadis et al. (2025), which were calculated by integrating the spectral energy distributions. The results are presented in Fig. 5 (we note here that the uncertainty in this estimate due to the correction is 0.05 dex; Neugent et al. 2020).
@@
From this plot, it is striking that there are several luminous sources above $5.5\,L_{\odot}$. In particular, we find 22 M31 and six M33 sources. In Table 3 we present them by including their Spitzer and Gaia DR3 IDs, their Gaia parallax and proper motions (whenever available), previous (original) and final (predicted) classifications, as well as their final probability, band completeness, and estimated luminosity.
@@
In M31, the three most luminous sources (close to $\textrm{log}(L/L_{\odot})\,$$\sim 7$; IDs: M31-439614, M31-439351, and M31-439254) do not have a Gaia counterpart, which means that these could be foreground red stars. However, the next most luminous source (ID M31-350) is a confirmed M1 I RSG (J004428.48+415130.9), with a range of $\textrm{log}(L/L_{\odot})\,$=5.435.64=5.43-5.64 McDonald et al. (2022). Such luminous sources may actually host a dense and complex circumstellar environment (similar to that observed in WOH G64; Ohnaka et al. 2024; Munoz-Sanchez et al. 2024b) that can contribute up to 0.3 dex in luminosity. If we consider these uncertainties in our approach, then our estimate of $\textrm{log}(L/L_{\odot})\,$=5.8=5.8 could be corrected by 0.35 dex, making it consistent with values from the literature. Based on our Gaia foreground cleaning method, the parallax and proper motion measurements make it a strong M31 candidate. Therefore, it certainly provides an upper bound for the HD limit in M31. Below that there are an additional 12 sources (between $\textrm{log}(L/L_{\odot})\,$$\sim 5.8-5.5$), of which three have Gaia parallax and proper motion values making them the best candidates of high luminous RSGs [^12] Among these sources is J003951.33+405303.7 from McDonald et al. (2022), corresponding to M31-439741, which we predicted robustly as an RSG. Sources J004731.12+422749.1 (M31-1011) and J004539.99+415404.1 (M31-336) were also predicted as RSG, but we were unable to estimate their luminosity. For sources J004428.12+415502.9 and J004520.67+414717.3 we were unable to find a good match within a few arcseconds in our catalog, probably due to our different Gaia cleaning approach., another set of eight sources with Gaia photometry only (with at least one band measurement), and one source without any Gaia match (hence, more likely to be a foreground source).
@@
In M33, we find two sources (M33-179 and M33-520) at $\textrm{log}(L/L_{\odot})\,$$\sim 6$). Since both of them lack Gaia parallax and proper motion measurements (but have photometric measurements), we cannot determine if they are genuine M33 members or not. These are followed by source M33-173 at $\textrm{log}(L/L_{\odot})\,$$\sim 5.7$, for which full Gaia information exists. However, this is a known OB star (Massey et al., 2016) and marks an erroneous prediction as an RSG. Another source (ID M33-646) has only Gaia photometry and, therefore, we can only tentatively consider it as a M33 RSG. The remaining two sources have full Gaia information and have been classified spectroscopically as RSGs. Therefore, they are the sources with the most robust data regarding their M31 membership status and spectral classification, and their derived luminosity sets a limit at $\textrm{log}(L/L_{\odot})\,$$\sim 6$ (which could be lowered to $\textrm{log}(L/L_{\odot})\,$$\sim 5.7$ if we consider the correction due to the circumstellar contribution).
[…]
@@ 4.4 Dusty YSGs
Figure 6: IR color ([3.6][4.5][3.6]-[4.5]) distribution of all (best candidate) sources predicted as YSGs. The violet curve corresponds to a Gaussian fit and the dashed black line indicates the $1\sigma$ threshold (0.24 mag) above which we identify sources as dusty YSGs.
Since our catalog is based on IR photometry, we are particularly sensitive to dusty, evolved stars. This makes us especially effective at detecting and characterizing the population of dusty YSGs, which are likely the evolutionary descendants of luminous RSGs. The ”RSG problem”, i.e., the observed scarcity of luminous RSGs exploding as Type-II supernovae (e.g., Smartt et al. 2009; Smartt 2015; Davies and Beasor 2020; Rodríguez 2022) suggests that the most luminous RSGs may evolve to warmer temperatures, either by undergoing a blue loop phase (e.g., Yoon and Cantiello 2010; Koumpia et al. 2020; Yang et al. 2023; Zapartas et al. 2025) or due to binary interaction because RSG winds are not strong enough to strip RSGs (e.g., Beasor et al. 2020; Decin et al. 2024; Antoniadis et al. 2024). Due to their high mass-loss rates during the RSG phase, these YSGs are often surrounded by significant amounts of material where dust actively forms (Gordon and Humphreys, 2019; Antoniadis et al., 2024; Decin et al., 2024).
To investigate potentially interesting candidates in our sample, we considered only the YSGs fulfilling the quality criteria (as set for final probability and band completeness in Sect. 2.4) in all galaxies. In Fig. 6 we show the distribution of their IR color [3.6][4.5][3.6]-[4.5]. We fit this distribution with a Gaussian function and found a mean value of -0.07 mag (consistent with 0, which is the expected mean value) and a standard deviation of 0.24 mag. This value is also consistent with the value (0.25 mag) that has been used previously (e.g., Kourniotis et al. 2017) based on the LMC and SMC results (see Bonanos et al. 2009, 2010). Hence, we find 159 sources above the $1\sigma$ threshold (indicated in Fig. 6), which we identify as dusty YSGs, corresponding to a fraction of $\sim 0.07$ with respect to their total number of 2071 sources. It is interesting to note that there are eight sources above the $2\sigma$ value (0.48 mag), namely M33-59527, M33-61917, M33-75218, M31-439225, M31-439242, M31-523223, WLM-19863, and NGC6822-30436), and another two (IC1613-14055 and WLM-234) above the $3\sigma$ threshold (0.72 mag). In Table 6 we present the IDs of all sources above the $1\sigma$ threshold. These are potentially yellow hypergiants that have suffered episodic mass loss and can provide a link to the ”RSG problem.” These sources need to be spectroscopically studied to verify their status.
5 Summary and conclusions
We investigated the effect of metallicity on different stellar populations, finding expected trends, such as a decrease in WR stars at lower metallicities and a relative increase in BSGs and YSGs. However, a number of selection biases, among which Spitzer ’s sensitivity to dusty evolved stars, must be considered when interpreting these trends. We also identified 21 luminous RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 5.5$), including six extreme RSGs in M31 ($\textrm{log}(L/L_{\odot})\,$$\geq 6$), challenging the HD limit. Further investigation of these sources is necessary to confirm their nature and more accurately determine their luminosity, as well as to handle carefully the statistical perplexity due to the small sample sizes. In addition, 159 dusty YSGs were detected. These are optimal candidates of yellow hypergiants, key sources to understand the ”RSG problem,” and follow-up observations are vital.
Our catalog provides the largest sample of machine-learning-classified massive stars in nearby galaxies, making it a crucial reference for future studies. It enables the identification of prime targets for spectroscopic follow-up, particularly luminous RSGs and YHGs, to further investigate their evolutionary pathways and the role of episodic mass loss in massive star evolution. Furthermore, we compiled a catalog of 5,273 spectroscopically confirmed sources for all 26 galaxies as derived from the literature. This offers a unique dataset for reference regarding spectral types for all massive stars and candidates known so far (including additional $\sim 330$ other sources), for galaxies beyond the Milky Way and the Clouds. The accuracy of our classifier can be enhanced by incorporating additional spectral data and improving the handling of observational biases. Expanding this method with JWST photometry will allow the study of massive stars in more distant galaxies, providing deeper insights into their role in galactic evolution.
Data availability
Tables 5 and 8 are only available in electronic form at CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
Acknowledgements.
The authors thank the referee for their useful comments. GM, AZB, KA, GMS, EC, SdW acknowledge funding support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (“ASSESS”, Grant agreement No. 772086). EZ acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project No: 7933). Facilities: This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The UHS is a partnership between the UK STFC, The University of Hawaii, The University of Arizona, Lockheed Martin and NASA. Software: This research made use of NumPy (numpy2020), pandas (pandas2010; pandas2011), matplotlib (matplotlib), imbalanced-learn (imbalanced-learn), scikit-learn (sklearn), SciPy (scipy2020), Astropy, a community-developed core Python package for Astronomy (astropy2013; astropy2018), Jupyter Notebooks (jupyter), TOPCAT, an interactive graphical viewer and editor for tabular data (topcat). The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of NASA’s Astrophysics Data System, and SIMBAD database, operated at CDS, Strasbourg, France.
Appendix A Detailed tables
In this section, we present a collection of tables regarding our data processing and results.
Table 4: Number of sources per photometric survey and foreground selection criteria per galaxy (see Sect. 2.3 for details.)  [^13] Spitzer data for galaxies IC 10, IC 1613, Pegasus DIG, Phoenix Dwarf, Sextans A and B, and WLM are derived from Boyer et al. (2015); M33, M81, NGC 2403, NGC 247, NGC 300, NGC 6822, and NGC 7793 from Khan et al. (2015); NGC 2366, NGC 253, NGC 4214, NGC 5253, and NGC 55 from Williams and Bonanos (2016); M31, M81, NGC 3077, NGC 4736, NGC 1313, and NGC 4826 from Khan (2017). \$*\$\$*\$footnotetext: For those galaxies the parallax and proper motion criteria derived from M31 were used, as there were not enough data to determine these quantities by using their data only.
[…]
Table 5: Final source catalog with predicted classifications for all galaxies.  [^14] This table is available in its entirety in VizieR/CDS catalog tool. A portion is shown here for guidance regarding its form and content.
ID RAJ2000 DEJ2000 Gaia_DR3_ID Final_Prob Final_Class Band_Compl
(deg) (deg)
WLM-1 0.52017 -15.44622 0.60202 WR 1.0
WLM-2 0.52012 -15.40681 0.64683 WR 0.4
WLM-3 0.52012 -15.36033 0.85906 RSG 1.0
WLM-4 0.52012 -15.46578 0.61820 WR 0.2
WLM-5 0.52012 -15.44331 0.64767 WR 0.2
WLM-6 0.52012 -15.41617 0.57403 WR 0.2
WLM-7 0.52008 -15.50714 0.61213 WR 0.2
WLM-8 0.52008 -15.46256 0.5356 WR 0.2
WLM-9 0.52008 -15.56622 0.43782 WR 0.6
WLM-10 0.52008 -15.42581 0.61508 WR 0.2
WLM-11 0.52008 -15.43997 0.61310 WR 0.2
WLM-12 0.52008 -15.54617 0.55765 WR 0.2
WLM-13 0.52008 -15.56514 0.39619 RSG 0.2
WLM-14 0.52004 -15.53664 0.47621 YSG 0.6
WLM-15 0.52004 -15.55456 0.47340 WR 0.2
WLM-16 0.52000 -15.40131 0.42985 WR 0.6
WLM-17 0.52000 -15.46006 0.64583 WR 0.2
WLM-18 0.52000 -15.60042 0.61512 WR 0.2
WLM-19 0.51996 -15.45511 0.61603 WR 0.2
WLM-20 0.51996 -15.35928 0.48667 RSG 0.2
@@ Appendix A Detailed tables
Sextans_A-608 NGC_2403-10650 M31-291254 IC_1613-2099 IC_1613-22555 NGC_6822-13118
M81-11303 M83-2819 M31-294149 IC_1613-2943 IC_1613-22602 NGC_6822-13324
NGC_253-1293 M83-4051 M31-347036 IC_1613-3830 IC_1613-23280 NGC_6822-13679
NGC_253-1478 M83-7927 M31-362279 IC_1613-4932 IC_1613-24144 NGC_6822-14299
NGC_253-1870 M83-13741 M31-380446 IC_1613-5603 IC_1613-24766 NGC_6822-14322
NGC_253-2676 Pegasus_DIG-6138 M31-409496 IC_1613-5966 IC_1613-25537 NGC_6822-14942
NGC_253-3150 Pegasus_DIG-8663 M31-409503 IC_1613-6832 WLM-69 NGC_6822-15128
M33-51717 Pegasus_DIG-8942 M31-409505 IC_1613-6921 WLM-179 NGC_6822-15582
M33-52312 M31-3384 M31-414106 IC_1613-7301 WLM-188 NGC_6822-16756
M33-53306 M31-$35$ M31-414142 IC_1613-8607 WLM-9147 NGC_6822-17168
M33-55564 M31-12298 M31-451363 IC_1613-10328 WLM-10681 NGC_6822-17404
M33-60369 M31-45371 M31-472507 IC_1613-10358 WLM-13121 NGC_6822-17503
M33-62164 M31-119542 M31-479549 IC_1613-10781 NGC_3109-2169 NGC_6822-17521
M33-62367 M31-129762 M31-484559 IC_1613-11142 NGC_3109-2807 NGC_6822-18213
M33-63872 M31-130114 M31-581217 IC_1613-11194 NGC_3109-5242 NGC_6822-19404
M33-65643 M31-132016 M31-714077 IC_1613-11517 NGC_6822-1590 NGC_6822-19748
M33-65743 M31-163495 M31-749160 IC_1613-12270 NGC_6822-2820 NGC_6822-20276
M33-65915 M31-173608 M31-775787 IC_1613-14279 NGC_6822-3299 NGC_6822-20731
M33-68313 M31-213432 M31-791111 IC_1613-16613 NGC_6822-5305 NGC_6822-21191
M33-69050 M31-217495 Sextans_B-1529 IC_1613-16662 NGC_6822-5378 NGC_6822-22201
M33-71044 M31-227660 Sextans_B-1691 IC_1613-17410 NGC_6822-5581 IC_10-289
M33-71771 M31-229975 Sextans_B-1745 IC_1613-19541 NGC_6822-7081 IC_10-858
NGC_4826-41 M31-246551 Sextans_B-3492 IC_1613-20037 NGC_6822-9860 IC_10-2758
NGC_4826-109 M31-256127 NGC_4736-847 IC_1613-20340 NGC_6822-10415 IC_10-19292
NGC_4826-3560 M31-261208 NGC_4736-4783 IC_1613-21452 NGC_6822-12447
NGC_2403-3659 M31-272080 IC_1613-221 IC_1613-21676 NGC_6822-12948
NGC_2403-5096 M31-273852 IC_1613-843 IC_1613-22503 NGC_6822-12966
Appendix B Classified sources from the literature